# Context is Key: A Benchmark for Forecasting with Essential Textual Information

### v1

@@ 5.4 Error Analysis
We find that models occasionally return forecasts that miss the ground truth by a large margin. A significant failure denotes a forecast that over or undershoots by at least five times the range of the ground truth; at that point, we clip the RCRPS to 5 as explained in Sec. 5.1. Despite this cap, such unpredictable behaviour impacts the results of Tab. 1: GPT-4o with Direct Prompt, while emerging as a top-performer in most tasks (as reflected in its average rank), provides significantly higher aggregate RCRPS than models ranked worse, such as Mixtral-8x7B with LLMP. As an example, Direct Prompt with GPT-4o fails significantly in a task with a context involving scientific notation (see Fig. 16; more examples can be found in Sec. C.5). These findings underscore the need for future work to develop more robust models that can handle context effectively while avoiding significant failures.
@@ 5.5 Inference Cost
A key practical aspect for forecasting applications is the inference time of models and their associated cost. Fig. 6 (left) shows that, while Llama-3.1-405B-Instruct has the best RCRPS, it comes at the cost of a significantly higher parameter count than the quantitative forecasters. This emphasizes that, while LLMs can be powerful context-aware forecasters, they come with a steep computational cost, highlighting the need for efficient models that balance both accuracy and resource demands. Of note is also that many LLM baselines are Pareto dominated by quantitative forecasters such as Lag-Llama and Chronos. This suggests that the ability to ingest text is not enough and that a careful choice of LLM and prompting strategy is crucial for Pareto efficiency.
Fig. 6 (right) emphasizes the disparity in inference time between LLMs and quantitative models. LLMs take significantly longer to make predictions, with the most accurate LLMs having inference times that are orders of magnitude higher than their quantitative counterparts. Quantitative models, in contrast, maintain much lower inference times, making them far more efficient for practical use. The high computational demands of context-aware LLMs hinder their practical use in real-world forecasting, especially where speed and cost matter. The clear benefits of incorporating context warrants research into making them more efficient, aiming to match the cost-effectiveness of traditional models and enabling their deployment in large-scale forecasting.
Figure 6: Overview of inference costs. (Left) Comparison of average RCRPS (per Tab. 1), vs. the parameter count of each baseline model (lower is better for both). The GPT family, as well as TimeGEN, are left out as there is no information on them about parameter count. The dashed line illustrates the Pareto front: models above and to the right of this front are dominated. Quantitative forecasters dominate the low-parameter regime, while LLM-based methods such as LLMP Llama-3-8B or 3-70B and DP 3.1-405B-Inst offer superior performance for a higher parameter count. (Right) Inference time in seconds, for all baselines, averaged over all tasks. Several quantitative methods are much faster on average than LLM-based methods. However, there are significant differences in inference time between the LLM-based forecasters: for the Llama models, LLM Process takes about an order of magnitude more time to run on average than Direct Prompt.
@@ 6 Related Work
We review two streams of related work: (i) work that introduce related benchmarks and datasets, and (ii) work that repurpose LLMs to obtain foundation models for context-aided forecasting.
@@
Benchmarks and Datasets  Merrill et al. (2024) present a benchmark designed to evaluate LLMs’ ability to reason about time series, with context-aided forecasting as one assessed capability. Their approach differs from ours in several important ways. First, they focus on purely synthetic time series, which may not accurately reflect real-world dynamics, whereas our benchmark is based primarily on real-world data. Second, their evaluation is limited to point forecasting metrics, which do not measure the quality of the full forecast distribution. In contrast, we adopt probabilistic forecasting metrics, e.g., the continuous ranked probability score (CRPS; c.f. Gneiting & Raftery2007), to assess the quality of entire forecast distributions. Other related datasets include Time-MMD (Liu et al., 2024a), which integrates text extracted from reports and web searches, TGTSF (Xu et al., 2024), which incorporates information such as weather reports and news articles, SysCaps (Emami et al., 2024), which includes LLM-generated descriptions of building energy consumption systems, and TS-Insights (Zhang et al., 2023), which includes LLM-generated descriptions of trends and seasonalities. The key distinction between these works and ours lies in the role of textual information: while in these works, the text is not essential to generating high-quality forecasts, in our benchmark, accurate forecasts cannot be achieved, by design, without effectively leveraging the provided textual information.
@@
Repurposing LLMs for Forecasting  A natural approach to this problem is to build forecasting methods based on LLMs. Xue & Salim (2023) showed that forecasting could be framed as a question-answering problem. Subsequently, Gruver et al. (2024) and Requeima et al. (2024) showed that zero-shot sequence completion could generate accurate forecasts and that textual side-information could be used to influence forecasts. However, their analysis is limited to illustrative examples rather than a comprehensive evaluation. Other approaches incorporate time series into pretrained LLMs (Jin et al., 2024; Liu et al., 2024b; Zhang et al., 2024) by introducing special tokens used to represent patched time series patterns; or modifying their encoders to account for time series data (Jia et al., 2024). While these methods show promising results, their evaluations primarily rely on datasets where the contextual information is not guaranteed to improve forecasts over numerical data alone. As a result, it remains unclear whether their success is driven by accurate numerical forecasting or by effectively incorporating context; this shortcoming motivates our investigation into this question.
@@ 7 Discussion
In this work, we propose the Context is Key (CiK) benchmark: a collection of forecasting tasks that require combining historical data with critical natural language context. We evaluate a range of models on CiK, including our proposed LLM prompting method, Direct Prompt, which achieves the best performance. We analyse and discuss the failure modes of these models, and our findings underscore the critical role of contextual information in improving forecasts, while also revealing both the unexpected strengths and notable limitations of the investigated LLM-based forecasters.
Limitations:  While our benchmark provides valuable insights into the integration of contextual information in time series forecasting, it is important to acknowledge its limitations. Our study excludes modalities other than time series data and text, and excludes multivariate time series scenarios. Although we carefully and deliberately designed the tasks to assess how well time series forecasters can integrate contextual information, our focus was on relationships between context and forecasts that are discernible to humans. Hence, our benchmark does not explicitly evaluate a models’ capacity to leverage latent relationships that might elude human observation. Moreover, while tasks are designed to require certain capabilities, we do not guarantee that alternative approaches to solving them do not exist. Our collection of capabilities and context types was not intended to be exhaustive but rather to serve as tools for analyzing forecasters’ performance on the benchmark. While we have taken steps to mitigate memorization concerns, as discussed in Sec. 3.1, achieving absolute certainty in this regard is challenging without strictly held-out data.
Future work:  There are several promising avenues for future work in the context-aided forecasting setting. Enhancements to the benchmark could include tasks that require multivariate forecasting or incorporate additional modalities, such as images, structured databases, or spatiotemporal data. Tasks that deliberately challenge context length limitations or probe specific weaknesses of language models would also be valuable additions. Furthermore, this benchmark strongly motivates research into developing more accurate and efficient multimodal forecasting models, which it is well-positioned to support. Lastly, as models become more robust, they could be integrated into agentic systems with conversational interfaces, allowing forecasts to be augmented with human expertise and automatically retrieved facts (e.g., via search engines). Such advancements would represent a significant step toward automating and democratizing access to powerful forecasting tools.
@@ Acknowledgements
The authors are grateful to Midan Kim, Torsten Scholak, Mohammad Reza Samsami, Oussama Boussif and Can Chen for their valuable feedback and suggestions. This research was supported by Mitacs Accelerate Grants and enabled by compute resources provided by ServiceNow Research and the Frontier supercomputer. The latter resources were awarded through the Frontier DD allocation and INCITE 2023 program for the project “Scalable Foundation Models for Transferable Generalist AI” and were supplied by the Oak Ridge Leadership Computing Facility at the Oak Ridge National Laboratory, with support from the Office of Science of the U.S. Department of Energy.
@@ A.1 Data Sources
A.2 Weighting scheme for tasks
To define the weight of each task, we first group the tasks in clusters. These clusters are primarily defined based on the original data source used to create the tasks. However, when tasks are fundamentally different, due to not testing the same capabilities, we put them in different clusters despite them using the same data source. For example, for tasks created using the Solar irradiance and cloud cover data, all of which ask models to forecast the irradiance, the tasks form three distinct clusters: one for tasks asking models to do forecast with very short history (less than a day), one for tasks giving the cloud cover as covariate, and the final one for tasks where the models are given a tight upper bound on the possible irradiance. Once we define these clusters, we simply equal weight to each cluster, and equal weight to each task inside each cluster.
A.3 Standard errors and average ranks
A.4 Model Capabilities
We provide a detailed explanation of each model capability here. Note that tasks in the CiK benchmark need not be mutually exclusive with the model capabilities they require; tasks are tagged with one or more model capabilities.
Instruction following (24 Tasks): Using direct instructions available in the context. Instructions could express constraints to be satisfied, or the expected effect of an event, for example.
@@ A.4 Model Capabilities
A.5 Task lengths
Fig. 7 provides an overview of the distribution of the lengths of the natural language context, numerical history and target (prediction horizon) for a set of five instances for each task in the CiK benchmark.
@@ Appendix B Examples of tasks from the benchmark
In this section, we feature multiple examples from the benchmark to exemplify exactly what a task is, what context sources represent (Sec. 3.2), and how these tasks encourage the use of capabilities (Sec. 3.3). To visualize all tasks in the benchmark, we refer the reader to https://servicenow.github.io/context-is-key-forecasting/v0/.

### v2

@@ 5.4 Error Analysis
Foundation models are known to make mistakes or return factually inaccurate information (Bommasani et al., 2021). We find that models occasionally return forecasts that miss the ground truth by a large margin. We use the term significant failure to denote a forecast that over or undershoots by at least $500\%$ of the range of the ground truth; we clip the RCRPS of such instances to 5 to avoid them disproportionately skewing the aggregate score. Despite this, such significant failures impact the results in Tab. 1: GPT-4o with direct prompt , while emerging as a top-performer in most tasks (as reflected in its average rank), has a significantly higher aggregate RCRPS than models ranked worse, such as Mixtral-8x7B with llmp. As an example, GPT-4o with direct prompt fails significantly in a task with a context involving scientific notation (see Figure 25 and more examples in Sec. C.12). Notably, while a model may generally achieve a high win rate, a few significant failures can dominate its aggregate performance, as observed for Mixtral-8x7B-Inst with direct prompt . We analyze this in detail in Sec. C.10. These findings underscore the need for future work to develop more robust models that can handle context effectively while avoiding significant failures.
@@ 5.5 Inference Cost
Figure 7: Comparison of average RCRPS (per Tab. 1), vs. the parameter count of each method(lower is better for both). The GPT family, as well as TimeGEN, are left out as there is no information on them about parameter count. The dashed line illustrates the Pareto front: models above and to the right of this line are dominated. Quantitative forecasters dominate the low-parameter regime, while LLM-based methods such as Qwen-2.5-7B-Inst ( direct prompt ) or Llama-3-70B (llmp) and 3.1-405B-Inst ( direct prompt ) offer superior performance for a higher parameter count.
A key practical aspect for forecasting applications is the inference time of models and their associated cost. Figure 7 shows that, while Llama-3.1-405B-Instruct performs the best on average, it comes at the cost of a significantly higher parameter count than the quantitative forecasters. This emphasizes that while LLMs can be powerful context-aware forecasters, they come with a steep computational cost, which highlights the need for efficient models that balance both accuracy and resource demands. Of note is also that many LLM-based forecasters are pareto-dominated quantitative forecasters such as Lag-Llama and Chronos . This suggests that, beyond the ability to simply process text, a careful choice of prompting strategy and LLM is crucial for Pareto efficiency. Further, due to their high parameter count, the LLMs have inference times that are orders of magnitude longer than quantitative forecasters which are far more efficient for sustained usage (see App. Figure 13 for a comparison). LLMs require significant computational power, making them unsuitable for real-world practical forecasting at scale where speed and cost matter. However, additional research could improve the efficiency of context-aided forecasting models to match the cost-effectiveness of traditional models.
@@ 6 Related Work
We review two streams of related work: (i) works that introduce related benchmarks and datasets, and (ii) works that repurpose LLMs for forecasting.
@@
Benchmarks and Datasets  Merrill et al. (2024) present a benchmark designed to evaluate LLMs’ ability to reason about time series, with context-aided forecasting as one assessed capability. They focus on purely synthetic time series, which may not accurately reflect real-world dynamics, whereas our benchmark is based primarily on real-world data. Further, their evaluation is limited to point forecasting metrics, which do not measure the quality of the full forecast distribution. In contrast, we adopt probabilistic forecasting metrics, e.g., the continuous ranked probability score (CRPS; c.f. Gneiting & Raftery2007), to assess the quality of entire forecast distributions. Other related datasets include Time-MMD (Liu et al., 2024a), which integrates text extracted from reports and web searches, TGTSF (Xu et al., 2024), which incorporates information such as weather reports and news articles, SysCaps (Emami et al., 2024), which includes LLM-generated descriptions of building energy consumption systems, TS-Insights (Zhang et al., 2023), which includes LLM-generated descriptions of trends and seasonalities, and Dual-Forecaster (Wu et al., 2025) where time series datasets are captioned with trend and seasonality information. Several works (Sawhney et al., 2021; Liu et al., 2024b; Wang et al., 2024b, a; Kim et al., 2024) automatically construct datasets of paired textual and numerical information. The key distinction between these works and ours lies in the importance of the textual information: while in the above works, the text is not guaranteed to be essential for high-quality forecasts, in CiK, all tasks are handcrafted to ensure that accurate forecasts cannot be achieved without relying on the provided natural language context, thereby making it a high-quality evaluation benchmark for context-aided forecasting.
@@
Repurposing LLMs for Forecasting  A natural approach to context-aided forecasting is to build methods based on LLMs. Xue & Salim (2023) showed that forecasting could be framed as a question-answering problem. Subsequently, Gruver et al. (2024) and Requeima et al. (2024) showed that LLMs could generate accurate forecasts with sequence completion, and that textual side-information could be used to influence forecasts. However, their analysis is limited to illustrative examples rather than a comprehensive evaluation. Other approaches have incorporated time series into pretrained LLMs (Jin et al., 2024; Liu et al., 2024c; Zhang et al., 2024) by introducing special tokens used to represent patched time series patterns, or modifying their encoders to account for time series data (Jia et al., 2024). While these methods show promising results, their evaluations primarily rely on datasets where the contextual information is not guaranteed to improve forecasts over numerical data alone. As a result, it remains unclear whether their performance is due by accurate numerical forecasting or by effectively incorporating context; this shortcoming motivates our investigation into this question.
@@ 7 Discussion
In this work, we propose the Context is Key (CiK) benchmark: a collection of forecasting tasks that require processing historical data with essential natural language context. We evaluate a range of models on CiK, including our proposed LLM prompting method, direct prompt , which achieves the best aggregate performance. We analyze and discuss the failure modes of these models, and our findings underscore the critical role of contextual information in improving forecasts, while also revealing both the unexpected strengths and notable limitations of the investigated LLM-based forecasters.
Limitations:  While our benchmark provides valuable insights into the integration of contextual information in time series forecasting, it is important to acknowledge its limitations. Our study limits itself to the natural language modality for context, and excludes multivariate time series scenarios. Although we deliberately designed the tasks to assess how well forecasting models can integrate contextual information, our benchmark does not evaluate whether models can leverage latent relationships that might elude human observation. While we have taken steps to mitigate memorization concerns, as discussed in Sec. 3.1, achieving absolute certainty in this regard is challenging without strictly held-out data.
Future work:  There are several promising avenues for future work. Extensions to CiK could include multivariate forecasting tasks, or tasks that incorporate other modalities such as images, databases or spatiotemporal data. Tasks that deliberately challenge context length limitations or probe specific weaknesses of language models would also be valuable additions. Methods to improve the automatic generation of large, high-quality datasets for context-aided forecasting could complement CiK. Furthermore, this benchmark strongly motivates research into developing more accurate and efficient multimodal forecasting models, which it is well-positioned to support. Lastly, as models become more robust, they could be integrated into agentic systems with conversational interfaces, allowing forecasts to be augmented with human expertise and automatically retrieved information. Such advancements would represent a significant step toward automating and democratizing access to powerful forecasting tools.
Impact Statement
This paper’s goal is to advance the development and refinement of forecasting methods which can leverage contextual information. It can thus increase the likelihood that such methods will be adopted by various organizations. Therefore, this paper shares the positive and negative impacts that such methods will have. Such methods could increase forecasting accuracy and efficiency, increasing the ability to foresee future courses and plan in consequence. A secondary effect is an increased democratization of access to high-quality forecasts. On the flip side, this democratization could increase reliance on automated methods for decision-making purposes.
@@ Acknowledgements
The authors are grateful to Andrei Dinin, Christian Hudon, Gabrielle Gauthier Melançon, Ghazwa Darwiche, Kiarash Mohammadi, Léo Boisvert, Megh Vipul Thakkar, Oluwanifemi Isaac Bamgbose, Orlando Marquez, Raymond Li, Thibault Le Sellier de Chezelles, Ethan Honey, Loic Mandine and Thomas Lai for participating in the human study on the relevance of context. The authors are grateful to Midan Kim, Torsten Scholak, Mohammad Reza Samsami, Oussama Boussif and Can Chen for their valuable feedback and suggestions. This research was supported by Mitacs Accelerate Grants and enabled by compute resources provided by ServiceNow Research and the Frontier supercomputer. The latter resources were awarded through the Frontier DD allocation and INCITE 2023 program for the project “Scalable Foundation Models for Transferable Generalist AI” and were supplied by the Oak Ridge Leadership Computing Facility at the Oak Ridge National Laboratory, with support from the Office of Science of the U.S. Department of Energy.
@@ A.1 Data Sources
A.2 Task Creation Process
All tasks were manually designed, from scratch, by the authors of this work without resorting to external annotators, crowdsourcing, or LLMs. We used the following procedure to create the tasks in the benchmark.
Task Diversity:  First, we identified high-quality sources of public time series data from various application domains (Sec. 3.1). Special care was taken to find data sources that are continuously updated to facilitate future benchmark updates. Second, we established the categorization for types of context (Sec. 3.2). Third, we posited various reasoning capabilities that one could potentially use to infer non-trivial information about numerical forecasts from contextual information (e.g., using common sense, making analogies to covariates, etc.; Sec. A.3). The task ideation process that followed aimed to ensure sufficient coverage of these three aspects.
Task Ideation:  With this framework established, all authors contributed to the ideation of new tasks. In summary, the process consisted of:
  1. 1.
    Selecting a data source
  2. 2.
    Implementing a time series window selection strategy (e.g., short or long history)
  3. 3.
    Brainstorming about the nature of contextual information that could help achieve better forecasts (e.g., information about the past) and the capabilities that might potentially serve to apply it to the forecasting problem.
  4. 4.
    Writing code to verbalize the context (e.g., calculating statistics of the series beyond the observed numerical history, creating a template to render such statistics as text, etc.), and
  5. 5.
    Finally, if required, writing code to modify the time series data to reflect the context (e.g., introducing some spikes in future values).
Peer Review:
Then, the tasks were peer-reviewed by a committee composed of all other authors (each with time series research experience). The creator of each task was not allowed to participate in the review. The review ensured that the contextual information was of high quality, that it undoubtedly enabled a better forecast, and that the context types used in the task were tagged correctly. If a task was deemed of not high enough quality, it was either returned for revisions or excluded from the benchmark.
Code availability:  The code for all tasks is available here: https://anonymous.4open.science/r/context-is-key-forecasting-E391/. An example task can be found here: https://anonymous.4open.science/r/context-is-key-forecasting-E391/cik_benchmark/tasks/montreal_fire/short_history.py, where the time series window selection occurs from L94-112 and the context generation occurs from L114-158.
A.3 Model Capabilities
As mentioned in Sec. A.2, we designed the tasks with consideration of the capabilities a model might potentially use to incorporate contextual information into its forecasts. All tasks in CiK are tagged with such capabilities. However, these tags are inherently subjective and not intended as formal attributions. Rather, they serve as broad categories to help readers identify examples of interest within the benchmark. These are as follows:
Common-Sense (24 Tasks): Using direct instructions available in the context. Instructions could express constraints to be satisfied, or the exact expected effect of an event, for example.
@@ A.3 Model Capabilities
[…]
A.4 Validating the Relevance of the Context
To evaluate the relevance of the contextual information for tasks in CiK, we query both human and LLM annotators on the relevance of the context. Our findings confirm that the context is relevant for all tasks (see Figure 8).
Figure 8: Ratings of the relevance of the context for both human annotations (x-axis) and LLM annotations (y-axis). There are 5 ratings per task for the LLM, and between 4 and 10 ratings per task for the human annotators. Each cell represents the number of tasks that correspond to a given pair of (human, LLM) ratings for the relevance of the context. For example, 53 (top right) of the 71 tasks have over 95% of annotators tagging the context as useful, as well as more than 95% of the LLM annotations tagging the context as useful. All tasks are considered relevant by either the LLM or the human annotators. Furthermore, the vast majority of tasks are considered relevant across more than 95% of ratings .
Figure 9: The task-wise proportion of annotators (n=11) who tag the context as useful. The overall rate of tasks tagged as useful across all annotators is 94.7%.
Figure 10: A histogram of results from the LLM-based critique of the relevance of context. Given the historical data, the future data and the associated context of tasks, GPT-4o is asked to assess whether its predictions would be “significantly better”, “slightly better”, “unchanged”, or “worse” (see Sec. A.4 for the details). The context in all tasks is considered as enabling better forecasts, with the majority of tasks having context that enable “significantly better” forecasts.
Human Evaluation of the Relevance of Context
To ensure that the context used in the tasks is relevant to the tasks, we ask 11 human annotators to evaluate the relevance of the context across 5 seeds. After a brief presentation of two example tasks based on examples from (Hyndman & Athanasopoulos, 2018), we ask the annotators whether the context enables better forecasts. Annotators are instructed that tasks are designed for the purpose of context-aided forecasting, and that we are seeking to identify tasks for which the context is not useful.
The results of this study can be found in Figure 9: the vast majority of tasks are always annotated as useful, while annotators disagree on a small minority of tasks, such as FullCausalContextImplicitEquationBivarLinSVAR (visualized at https://anon-forecast.github.io/benchmark_report_dev/FullCausalContextImplicitEquationBivarLinSVAR.html) likely due to the highly statistical nature of such task(s). Overall, $94.7\%$ of annotations report that the context is useful.
An LLM-based Critique of the Relevance of Context
To further validate the quality of the tasks, we build an LLM-based critique by prompting GPT-4o with the historical and future numerical data, as well as the context, and asking it to assess whether its estimation of future values would be “significantly better”, “slightly better”, “unchanged”, or “worse” when the context is provided compared to when it is not provided. Note that this experiment was ran after the benchmark was created, as an analysis tool to further validate the quality of the tasks.
We run this critique on 5 instances of each of the 71 tasks and report results in Figure 10222For Figure 8, we consider both “significantly better” and “slightly better” as meaning the context is useful. All tasks are assessed as enabling better forecasts when given context, with the majority of tasks assessed as having contexts that enable “significantly better” forecasts. The code linked to this experiment is provided at https://github.com/anon-forecast/benchmark_report_dev/blob/main/iclr_rebuttal_resources/llm_validation.py. The prompt used in the critique is in Figure 11.
You are a critic whose role is to evaluate the quality of tasks in the "context is key" time series forecasting
benchmark.
"Context is Key" (CiK) is a time series forecasting benchmark that pairs numerical data with diverse types of
carefully crafted textual context, requiring models to integrate both modalities to arrive at accurate
predictions.
Here is a task to evaluate.
<history>
((history))
</history>
<context>
<background>
((background))
</background>
<scenario>
((scenario))
</scenario>
<constraints>
((constraints))
</constraints>
</context>
<future>
((future))
</future>
Assume the following two scenarios:
1) You are given only the numerical data in <history> and have no additional information about the nature of the
time series. You must ignore the <context> section completely.
2) You are given the <context> section in addition to the numerical data in <history>.
Now, assume you had to estimate the probability distribution of the <future> values given the information
available in each scenario. How would the quality of your estimation change in scenario 2 compared to
scenario 1?
First show your reasoning in <reason></reason> tags, then answer in <answer></answer> tags with either
"significantly better", "slightly better", "unchanged", "worse" (no other responses are allowed).
Figure 11: The prompt used to query the LLM critique.
A.5 Weighting scheme for tasks
To define the weight of each task, we first group the tasks in clusters. These clusters are primarily defined based on the original data source used to create the tasks. However, when tasks are fundamentally different, due to not testing the same capabilities, we put them in different clusters despite them using the same data source. For example, for tasks created using the Solar irradiance and cloud cover data, all of which ask models to forecast the irradiance, the tasks form three distinct clusters: one for tasks asking models to do forecast with very short history (less than a day), one for tasks giving the cloud cover as covariate, and the final one for tasks where the models are given a tight upper bound on the possible irradiance. Once we define these clusters, we simply give equal weight to each cluster, and equal weight to each task inside each cluster.
A.6 Standard errors and average ranks
A.7 Task lengths
@@
Figure 12 provides an overview of the distribution of the lengths of the natural language context, numerical history and target (prediction horizon) for a set of five instances for each task in the CiK benchmark.
@@ Appendix B Examples of tasks from the benchmark
In this section, we feature multiple examples from the benchmark to exemplify exactly what a task is, what context types represent (Sec. 3.2), and how we tag these tasks with descriptive capabilities (Sec. A.3). To visualize all tasks in the benchmark, we refer the reader to https://anon-forecast.github.io/benchmark_report_dev.